! Required information [The following information applies to the questions displayed below.) Vernon Company is a retail compaUse this information to prepare a selling and administrative expenses budget. f. Utilities and sales commissions are paid theUse this information to prepare a selling and administrative expenses budget. f. Utilities and sales commissions are paid thef. Utilities and sales commissions are paid the month after they are incurred; all other expenses are paid in the month in whf. Utilities and sales commissions are paid the month after they are incurred; all other expenses are paid in the month in whf. Utilities and sales commissions are paid the month after they are incurred; all other expenses are paid in the month in whf. Utilities and sales commissions are paid the month after they are incurred; all other expenses are paid in the month in whRequired A Required B Required c Required D Required E Required F Required G Vernon borrows funds, in increments of $1,000, a
Answer #1
a) Sales budget
October November December
Cash Sales $    126,000.00 $  157,500.00 $    196,875.00
Sales on Account $    154,000.00 $  192,500.00 $    240,625.00
Total Budgeted Sales $    280,000.00 $  350,000.00 $    437,500.00
b) Schedule of cash receipts
October November December
Current Cash sales $    126,000.00 $  157,500.00 $    196,875.00
Add: Collection From A/R $                   -    $  154,000.00 $    192,500.00
Total Collections $    126,000.00 $  311,500.00 $    389,375.00
c)  Inventory purchase budget
October November December
Budgeted Cost of goods sold $    168,000.00 $  210,000.00 $    262,500.00
Add: Desired Ending Inventory $      21,000.00 $    26,250.00 $      13,600.00
Inventory Needed $    189,000.00 $  236,250.00 $    276,100.00
Less: Beginning Inventory $      16,800.00 $    21,000.00 $      26,250.00
Required Purchases on account $    172,200.00 $  215,250.00 $    249,850.00
d. Schedule of Cash payments budget for Inventory Purchases
October November December
Payment of Current Month's of  accounts payable $    137,760.00 $  172,200.00 $    199,880.00
Payment of Prior  Month's of  accounts payable $    34,440.00 $      43,050.00
Total Budgeted payment of Inventory $    137,760.00 $  206,640.00 $    242,930.00
e. Selling & Administrative Exp budget
October November December
Salary (fixed) $      19,600.00 $    19,600.00 $      19,600.00
Sales commissions (5% of sales) $    14,000.00 $      17,500.00
Supplies expense (2% of sales) $        5,600.00 $      7,000.00 $        8,750.00
Utilities (fixed) $        3,000.00 $      3,000.00 $        3,000.00
Depreciation $        5,600.00 $      5,600.00 $        5,600.00
Rent (fixed) $        6,400.00 $      6,400.00 $        6,400.00
Miscellaneous (fixed) $        2,800.00 $      2,800.00 $        2,800.00
Total $      43,000.00 $    58,400.00 $      63,650.00

A B С D October =B24-45% =B24-55% 280000 November =C24*45% =C2455% =B24.1.25 December =D24-45% =D24*55% =C24.1.25 October =B2